# How Uniform Random Weights Induce Non-uniform Bias: Typical Interpolating Neural Networks Generalize with Narrow Teachers

### v1

@@ B.2 Neuron-Scaled Fully-Connected Neural Networks
so using the notation from Def. 5.1 we find that
$\bm{\gamma}^{\left(1\right)}\odot\mathbf{W}^{\left(1\right)}f_{D}^{\left(0\right)}\left(\mathbf{x}\right)+\mathbf{b}^{\left(1\right)}=\bm{\gamma}^{\left(1\right)}\odot\mathbf{W}^{\left(1\right)}\mathbf{x}+\mathbf{b}^{\left(1\right)}=\left[\begin{array}[]{c}\bm{\gamma}^{\left(1\right)}_{1}\odot\mathbf{W}_{11}^{\left(1\right)}\mathbf{x}+\mathbf{b}_{1}^{\left(1\right)}\\ \bm{\gamma}^{\left(1\right)}_{2}\odot\mathbf{W}_{21}^{\left(1\right)}\mathbf{x}+\mathbf{b}_{2}^{\left(1\right)}\end{array}\right]$
@@
$f_{D}^{\left(1\right)}\left(\mathbf{x}\right)=\sigma\left(\left[\begin{array}[]{c}\bm{\gamma}^{\left(1\right)}_{1}\odot\mathbf{W}_{11}^{\left(1\right)}\mathbf{x}+\mathbf{b}_{1}^{\left(1\right)}\\ \bm{\gamma}^{\left(1\right)}_{2}\odot\mathbf{W}_{21}^{\left(1\right)}\mathbf{x}+\mathbf{b}_{2}^{\left(1\right)}\end{array}\right]\right)=\left[\begin{array}[]{c}\sigma\left(\bm{\gamma}^{\left(1\right)}_{1}\odot\mathbf{W}_{11}^{\left(1\right)}\mathbf{x}+\mathbf{b}_{1}^{\left(1\right)}\right)\\ \sigma\left(\bm{\gamma}^{\left(1\right)}_{2}\odot\mathbf{W}_{21}^{\left(1\right)}\mathbf{x}+\mathbf{b}_{2}^{\left(1\right)}\right)\end{array}\right]\,.$
@@
$\pi_{1}\left(f_{D}^{\left(1\right)}\left(\mathbf{x}\right)\right)=\sigma\left(\bm{\gamma}^{\left(1\right)}_{1}\odot\mathbf{W}_{11}^{\left(1\right)}\mathbf{x}+\mathbf{b}_{1}^{\left(1\right)}\right)\,.$ (35)
Since $h_{\bm{\theta}}\in\mathcal{E}$, we have that $\mathbf{W}_{11}^{\left(1\right)}=\mathbf{W}^{\star\left(1\right)}$, $\mathbf{b}_{1}^{\left(1\right)}=\mathbf{b}^{\star\left(1\right)}$ and $\bm{\gamma}_{1}^{\left(1\right)}=\bm{\gamma}^{\star\left(1\right)}$, so from (35) the coordinate projection is
$\pi_{1}\left(f_{D}^{\left(1\right)}\left(\mathbf{x}\right)\right)=\sigma\left(\bm{\gamma}^{\star\left(1\right)}\odot\mathbf{W}^{\star\left(1\right)}\mathbf{x}+\mathbf{b}^{\star\left(1\right)}\right)=f_{D^{\star}}^{\left(1\right)}\left(\mathbf{x}\right)\,.$ (36)
In addition, since $\bm{\gamma}_{2}^{\left(l\right)}=\mathbf{0}_{d_{l}-d_{l}^{\star}}$ and also $\mathbf{b}_{2}^{\left(1\right)}=\mathbf{0}_{d_{l}-d_{l}^{\star}}$, as well as $\sigma\left(0\right)=0$, we have that
$\sigma\left(\bm{\gamma}^{\left(1\right)}_{2}\odot\mathbf{W}_{21}^{\left(1\right)}\mathbf{x}+\mathbf{b}_{2}^{\left(1\right)}\right)=\mathbf{0}_{d_{l}-d_{l}^{\star}}\,.$ (37)
Putting 37 and 36 together we have
@@
$=\sigma\left(\bm{\gamma}^{\left(l\right)}\odot\mathbf{W}^{\left(l\right)}f_{D}^{\left(l-1\right)}\left(\mathbf{x}\right)+\mathbf{b}^{\left(l\right)}\right)$
$=\sigma\left(\bm{\gamma}^{\left(l\right)}\odot\left[\begin{array}[]{cc}\mathbf{W}_{11}^{\left(l\right)}&\mathbf{W}_{12}^{\left(l\right)}\\ \mathbf{W}_{21}^{\left(l\right)}&\mathbf{W}_{22}^{\left(l\right)}\end{array}\right]\left[\begin{array}[]{c}f_{D^{\star}}^{\left(l-1\right)}\left(\mathbf{x}\right)\\ \mathbf{0}_{d_{l-1}-d^{\star}_{l-1}}\end{array}\right]+\left[\begin{array}[]{c}\mathbf{b}_{1}^{\left(l\right)}\\ \mathbf{b}_{2}^{\left(l\right)}\end{array}\right]\right)$
$=\sigma\left(\bm{\gamma}^{\left(l\right)}\odot\left[\begin{array}[]{c}\mathbf{W}_{11}^{\left(l\right)}\\ \mathbf{W}_{21}^{\left(l\right)}\end{array}\right]f_{D^{\star}}^{\left(l-1\right)}\left(\mathbf{x}\right)+\left[\begin{array}[]{c}\mathbf{b}_{1}^{\left(l\right)}\\ \mathbf{b}_{2}^{\left(l\right)}\end{array}\right]\right)$
$=\sigma\left(\left[\begin{array}[]{c}\bm{\gamma}_{1}^{\left(l\right)}\odot\mathbf{W}_{11}^{\left(l\right)}f_{D^{\star}}^{\left(l-1\right)}\left(\mathbf{x}\right)+\mathbf{b}_{1}^{\left(l\right)}\\ \bm{\gamma}_{2}^{\left(l\right)}\odot\mathbf{W}_{21}^{\left(l\right)}f_{D^{\star}}^{\left(l-1\right)}\left(\mathbf{x}\right)+\mathbf{b}_{2}^{\left(l\right)}\end{array}\right]\right)$
$=\left[\begin{array}[]{c}\sigma\left(\bm{\gamma}_{1}^{\left(l\right)}\odot\mathbf{W}_{11}^{\left(l\right)}f_{D^{\star}}^{\left(l-1\right)}\left(\mathbf{x}\right)+\mathbf{b}_{1}^{\left(l\right)}\right)\\ \sigma\left(\bm{\gamma}_{2}^{\left(l\right)}\odot\mathbf{W}_{21}^{\left(l\right)}f_{D^{\star}}^{\left(l-1\right)}\left(\mathbf{x}\right)+\mathbf{b}_{2}^{\left(l\right)}\right)\end{array}\right]$
$=\left[\begin{array}[]{c}\sigma\left(\bm{\gamma}^{\star\left(l\right)}\odot\mathbf{W}^{\star\left(l\right)}f_{D^{\star}}^{\left(l-1\right)}\left(\mathbf{x}\right)+\mathbf{b}^{\star\left(l\right)}\right)\\ \mathbf{0}_{d_{l-1}-d^{\star}_{l-1}}\end{array}\right]$
@@
For the last layer, $l=L$, the proof is identical, except for the scalar product and the application of the activation function, which are removed. Since we assume that $h_{\bm{\theta}}$ and $h_{\bm{\theta}^{\star}}$ have the same output dimension $d_{L}$, this proves that for all $\mathbf{x}\in\mathbb{R}^{d_{0}}$
$h_{\bm{\theta}}\left(\mathbf{x}\right)=\pi_{L}\left(h_{\bm{\theta}}\left(\mathbf{x}\right)\right)=h_{\bm{\theta}^{\star}}\left(\mathbf{x}\right)\,.$
That is, $\mathcal{E}\subseteq\left\{h_{\bm{\theta}}\middle|h_{\bm{\theta}}\equiv h_{\bm{\theta}^{\star}}\right\}$ and therefore
$\mathbb{P}\left(\mathcal{E}\right)\leq\mathbb{P}\left(h_{\bm{\theta}}\equiv h_{\bm{\theta}^{\star}}\right)\,.$
Finally, to calculate the probability to sample $h_{\bm{\theta}}$ in $\mathcal{E}$ we count the number of constrained parameters - parameters which are either determined by $h_{\bm{\theta}^{\star}}$ or are 00 in $\mathcal{E}$. Looking at the dimensions of $\mathbf{W}_{11}^{\left(l\right)}$, $\mathbf{b}_{1}^{\left(l\right)}$, $\mathbf{b}_{2}^{\left(l\right)}$ and $\bm{\gamma}^{\left(l\right)}$, we deduce that there are exactly
@@
As in the proof of Theorem B.3, under the uniform prior over parameters $\mathcal{P}$,
@@
B.3 Convolutional Neural Network
@@
Definition B.7 (CNN restated).
For multi-channel inputs $\mathbf{x}^{j},\,\forall j\in\left[c_{0}\right]$, depth $L$, activation function $\sigma\left(\cdot\right)$, and multi-index channels number $D=\left(c_{1},\dots,c_{L},d_{s}\right)$, $\mathbf{h}^{\mathrm{CNN}}_{\bm{\theta}}$ is a convolutional network (CNN) defined recursively, starting with $\forall j\in\left[c_{0}\right]$, $\mathbf{f}_{\mathrm{CNN}}^{\left(0\right),j}\left(\mathbf{x}\right)=\mathbf{x}^{j}$, and then, for all $l\in\left[L\right]$ and $i\in\left[c_{l}\right]$, as
@@
$\mathbf{h}^{\mathrm{CNN}}_{\bm{\theta}}\left(\mathbf{x}\right)=$ $\operatorname{sign}\left(\mathrm{Vec}\left(\mathbf{f}_{\mathrm{CNN}}^{\left(L-1\right)}\left(\mathbf{x}\right)\right)^{\top}\mathbf{w}^{\left(L+1\right)}+b^{\left(L+1\right)}\right)\,,$
@@
$\mathbf{\bm{\theta}}=\left(\left\{\mathbf{K}^{\left(l\right)}\right\}_{l=1}^{L},\left\{\mathbf{b}^{\left(l\right)}\right\}_{l=1}^{L},\mathbf{w}^{\left(L+1\right)},b^{\left(L+1\right)}\right)\,,$
@@
Motivation. In order to tackle the analysis of a convolutional neural network at ease, we will define some new operations. We observe that our previous arguments from Section B.1 is agnostic to the convolution itself and relies only on parameter counting. In general, CNNs are analogous to FCNs in the sense that with spatial dimension of $1$, a convolutional layer is equivalent to a layer in a FCN, with each channel analogous to a neuron. Therefore, we define:
Definition B.9.
@@
and Def. B.9 we can write the multi-channel convolution in block form as
@@
Specifically, we can represent a convolutional layer, as defined in Def. B.7 by
@@
Theorem B.10 (Teacher equivalence probability for CNN restated).
For any activation function and any depth $L$, let $D^{\star}=\left(c_{1}^{\star},\dots,c_{L}^{\star},d_{s}^{\star}\right),\;D=\left(c_{1},\dots,c_{L},d_{s}\right)\in\mathbb{N}^{L+1}$ such that $D^{\star}\leq D$. If there exists some teacher $h_{\bm{\theta}^{\star}}\in\mathcal{H}^{\mathrm{CNN}}_{D^{\star}}$ and the prior is $\mathcal{P}=\mathcal{P}\left(\mathcal{H}^{\mathrm{CNN}}_{D}\right)$ then
@@
And by Theorem 3.2, $N=(\tilde{C}+3\log 2/\delta)/\epsilon$ samples are enough to ensure that for posterior sampling (i.e. G&C), $\mathcal{L}(\mathcal{A}_{\mathcal{P}}(\mathcal{S}))\leq\epsilon$ with probability $1-\delta$ over $\mathcal{S}\sim\mathcal{D}^{N}$ and the sampling.
@@
We follow the same strategy as in the proof of Theorem B.3. We define a sufficient condition on the parameters of the student network to ensure teacher equivalence, and then find the probability that this condition holds. Recall that for all $l\in\left[L\right]$, $\mathbf{K}^{\star\left(l\right)}$ and $\mathbf{b}^{\star\left(l\right)}$, where $\mathbf{K}_{i,j}^{\star\left(l\right)}\in\mathbb{R}^{k_{l}}$ is a convolution kernel with $k_{l}$ parameters, and $\mathbf{b}^{\star\left(l\right)}\in\mathbb{R}^{c_{l}}$, are the teacher’s $l^{\text{th}}$ layer’s convolution kernels and bias terms, respectively. Using the notation introduced in Def. B.9, the teacher’s convolution kernels can be arranged as
@@
=$\mathcal{E}=\left\{h_{\bm{\theta}}\in\mathcal{H}_{D}^{\text{CNN}}\middle|\begin{array}[]{l}\forall l=1,\dots,L\;\mathbf{K}^{\left(l\right),\left(11\right)}=\mathbf{K}^{\star\left(l\right)},\;\mathbf{K}^{\left(l\right),\left(12\right)}=\mathbf{0},\;\mathbf{b}_{1}^{\left(l\right)}=\mathbf{b}^{\star\left(l\right)},\\ \mathbf{w}^{\left(L+1\right),1}=\mathbf{w}^{\star\left(L+1\right)},\;\mathbf{w}^{\left(L+1\right),2}=\mathbf{0}_{d_{s}-d_{s}^{\star}},\;b^{\left(L+1\right)}=b^{\star\left(L+1\right)}\end{array}\right\}\,.$
We claim that any $h_{\bm{\theta}}\in\mathcal{E}$ is TE and can show this by induction over the layer $l$. With block notation, it is easy to see that this is identical to the proof of Theorem B.3 when substituting $\mathbf{W}^{\left(l\right)}$ with $\mathbf{K}^{\left(l\right)}$, $f_{\mathrm{FC}}^{\left(l\right)}$ with $\mathbf{f}_{\mathrm{CNN}}^{\left(l\right)}$, and standard matrix multiplication with convolution. Finally, we turn to finding $\mathbb{P}\left(\mathcal{E}\right)$. For a CNN $\mathbf{h}^{\mathrm{CNN}}_{\bm{\theta}}\in\mathcal{E}$, there are $c_{l}^{\star}\cdot c_{l-1}^{\star}$ convolution kernels set to equal the teacher’s, $c_{l}^{\star}\cdot\left(c_{l-1}-c_{l-1}^{\star}\right)$ kernels set to 00, and $c_{l}^{\star}$ bias terms. Each convolution kernel $\mathbf{K}_{i,j}^{\left(l\right)}$ is defined by $k_{l}$ parameters. From the linear layer, we need to account for $d_{s}$ weight parameters and one bias term. In total,
@@
B.4 Channel-Scaled Convolutional Neural Networks
Analogous to the FCN case, we present an additional CNN architecture. For simplicity, we state the definition of the architecture using the previously defined notation from B.9 and (39).
Definition B.11 (Channel Scaled CNN).
For multi-channel inputs $\mathbf{x}^{j},\,\forall j\in\left[c_{0}\right]$, depth $L$, activation function $\sigma\left(\cdot\right)$, and multi-index channels number $D=\left(c_{1},\dots,c_{L},d_{s}\right)$, $\mathbf{h}^{\mathrm{SCN}}_{\bm{\theta}}$ is a channel scaled convolutional neural network, or scaled convolutional network (SCN) for short, defined recursively, starting with $\mathbf{f}_{\mathrm{SCN}}^{\left(0\right)}\left(\mathbf{x}\right)=\mathbf{x}$, and then, for all $l\in\left[L\right]$, as
$\mathbf{f}_{\mathrm{SCN}}^{\left(l\right)}\left(\mathbf{x}\right)=$ $\left(\bm{\gamma}^{\left(l\right)}\odot\sigma\mathbf{K}^{\left(l\right)}*\mathbf{f}_{\mathrm{SCN}}^{\left(l-1\right)}+\mathbf{b}^{\left(l\right)}\right)$ (40)
$\mathbf{h}^{\mathrm{SCN}}_{\bm{\theta}}\left(\mathbf{x}\right)=$ $\operatorname{sign}\left(\mathrm{Vec}\left(\mathbf{f}_{\mathrm{SCN}}^{\left(L\right)}\left(\mathbf{x}\right)\right)^{\top}\mathbf{w}^{\left(L+1\right)}+b^{\left(L+1\right)}\right)\,,$

### v2

@@ C.2 Neuron-Scaled Fully-Connected Neural Networks
so using the notation from Def. 4.1 we find that
$\boldsymbol{\gamma}^{\left(1\right)}\odot\mathbf{W}^{\left(1\right)}f_{D}^{\left(0\right)}\left(\mathbf{x}\right)+\mathbf{b}^{\left(1\right)}=\boldsymbol{\gamma}^{\left(1\right)}\odot\mathbf{W}^{\left(1\right)}\mathbf{x}+\mathbf{b}^{\left(1\right)}=\left[\begin{array}[]{c}\boldsymbol{\gamma}^{\left(1\right)}_{1}\odot\mathbf{W}_{11}^{\left(1\right)}\mathbf{x}+\mathbf{b}_{1}^{\left(1\right)}\\ \boldsymbol{\gamma}^{\left(1\right)}_{2}\odot\mathbf{W}_{21}^{\left(1\right)}\mathbf{x}+\mathbf{b}_{2}^{\left(1\right)}\end{array}\right]$
@@
$f_{D}^{\left(1\right)}\left(\mathbf{x}\right)=\sigma\left(\left[\begin{array}[]{c}\boldsymbol{\gamma}^{\left(1\right)}_{1}\odot\mathbf{W}_{11}^{\left(1\right)}\mathbf{x}+\mathbf{b}_{1}^{\left(1\right)}\\ \boldsymbol{\gamma}^{\left(1\right)}_{2}\odot\mathbf{W}_{21}^{\left(1\right)}\mathbf{x}+\mathbf{b}_{2}^{\left(1\right)}\end{array}\right]\right)=\left[\begin{array}[]{c}\sigma\left(\boldsymbol{\gamma}^{\left(1\right)}_{1}\odot\mathbf{W}_{11}^{\left(1\right)}\mathbf{x}+\mathbf{b}_{1}^{\left(1\right)}\right)\\ \sigma\left(\boldsymbol{\gamma}^{\left(1\right)}_{2}\odot\mathbf{W}_{21}^{\left(1\right)}\mathbf{x}+\mathbf{b}_{2}^{\left(1\right)}\right)\end{array}\right]\,.$
@@
$\pi_{1}\left(f_{D}^{\left(1\right)}\left(\mathbf{x}\right)\right)=\sigma\left(\boldsymbol{\gamma}^{\left(1\right)}_{1}\odot\mathbf{W}_{11}^{\left(1\right)}\mathbf{x}+\mathbf{b}_{1}^{\left(1\right)}\right)\,.$ (36)
Since $h_{\boldsymbol{\theta}}\in\mathcal{E}$, we have that $\mathbf{W}_{11}^{\left(1\right)}=\mathbf{W}^{\star\left(1\right)}$, $\mathbf{b}_{1}^{\left(1\right)}=\mathbf{b}^{\star\left(1\right)}$ and $\boldsymbol{\gamma}_{1}^{\left(1\right)}=\boldsymbol{\gamma}^{\star\left(1\right)}$, so from (36) the coordinate projection is
$\pi_{1}\left(f_{D}^{\left(1\right)}\left(\mathbf{x}\right)\right)=\sigma\left(\boldsymbol{\gamma}^{\star\left(1\right)}\odot\mathbf{W}^{\star\left(1\right)}\mathbf{x}+\mathbf{b}^{\star\left(1\right)}\right)=f_{D^{\star}}^{\left(1\right)}\left(\mathbf{x}\right)\,.$ (37)
In addition, since $\boldsymbol{\gamma}_{2}^{\left(l\right)}=\mathbf{0}_{d_{l}-d_{l}^{\star}}$ and also $\mathbf{b}_{2}^{\left(1\right)}=\mathbf{0}_{d_{l}-d_{l}^{\star}}$, as well as $\sigma\left(0\right)=0$, we have that
$\sigma\left(\boldsymbol{\gamma}^{\left(1\right)}_{2}\odot\mathbf{W}_{21}^{\left(1\right)}\mathbf{x}+\mathbf{b}_{2}^{\left(1\right)}\right)=\mathbf{0}_{d_{l}-d_{l}^{\star}}\,.$ (38)
Putting 38 and 37 together we have
@@
$=\sigma\left(\boldsymbol{\gamma}^{\left(l\right)}\odot\mathbf{W}^{\left(l\right)}f_{D}^{\left(l-1\right)}\left(\mathbf{x}\right)+\mathbf{b}^{\left(l\right)}\right)$
$=\sigma\left(\boldsymbol{\gamma}^{\left(l\right)}\odot\left[\begin{array}[]{cc}\mathbf{W}_{11}^{\left(l\right)}&\mathbf{W}_{12}^{\left(l\right)}\\ \mathbf{W}_{21}^{\left(l\right)}&\mathbf{W}_{22}^{\left(l\right)}\end{array}\right]\left[\begin{array}[]{c}f_{D^{\star}}^{\left(l-1\right)}\left(\mathbf{x}\right)\\ \mathbf{0}_{d_{l-1}-d^{\star}_{l-1}}\end{array}\right]+\left[\begin{array}[]{c}\mathbf{b}_{1}^{\left(l\right)}\\ \mathbf{b}_{2}^{\left(l\right)}\end{array}\right]\right)$
$=\sigma\left(\boldsymbol{\gamma}^{\left(l\right)}\odot\left[\begin{array}[]{c}\mathbf{W}_{11}^{\left(l\right)}\\ \mathbf{W}_{21}^{\left(l\right)}\end{array}\right]f_{D^{\star}}^{\left(l-1\right)}\left(\mathbf{x}\right)+\left[\begin{array}[]{c}\mathbf{b}_{1}^{\left(l\right)}\\ \mathbf{b}_{2}^{\left(l\right)}\end{array}\right]\right)$
$=\sigma\left(\left[\begin{array}[]{c}\boldsymbol{\gamma}_{1}^{\left(l\right)}\odot\mathbf{W}_{11}^{\left(l\right)}f_{D^{\star}}^{\left(l-1\right)}\left(\mathbf{x}\right)+\mathbf{b}_{1}^{\left(l\right)}\\ \boldsymbol{\gamma}_{2}^{\left(l\right)}\odot\mathbf{W}_{21}^{\left(l\right)}f_{D^{\star}}^{\left(l-1\right)}\left(\mathbf{x}\right)+\mathbf{b}_{2}^{\left(l\right)}\end{array}\right]\right)$
$=\left[\begin{array}[]{c}\sigma\left(\boldsymbol{\gamma}_{1}^{\left(l\right)}\odot\mathbf{W}_{11}^{\left(l\right)}f_{D^{\star}}^{\left(l-1\right)}\left(\mathbf{x}\right)+\mathbf{b}_{1}^{\left(l\right)}\right)\\ \sigma\left(\boldsymbol{\gamma}_{2}^{\left(l\right)}\odot\mathbf{W}_{21}^{\left(l\right)}f_{D^{\star}}^{\left(l-1\right)}\left(\mathbf{x}\right)+\mathbf{b}_{2}^{\left(l\right)}\right)\end{array}\right]$
$=\left[\begin{array}[]{c}\sigma\left(\boldsymbol{\gamma}^{\star\left(l\right)}\odot\mathbf{W}^{\star\left(l\right)}f_{D^{\star}}^{\left(l-1\right)}\left(\mathbf{x}\right)+\mathbf{b}^{\star\left(l\right)}\right)\\ \mathbf{0}_{d_{l-1}-d^{\star}_{l-1}}\end{array}\right]$
@@
For the last layer, $l=L$, the proof is identical, except for the scalar product and the application of the activation function, which are removed. Since we assume that $h_{\boldsymbol{\theta}}$ and $h_{\boldsymbol{\theta}^{\star}}$ have the same output dimension $d_{L}$, this proves that for all $\mathbf{x}\in\mathbb{R}^{d_{0}}$
$h_{\boldsymbol{\theta}}\left(\mathbf{x}\right)=\pi_{L}\left(h_{\boldsymbol{\theta}}\left(\mathbf{x}\right)\right)=h_{\boldsymbol{\theta}^{\star}}\left(\mathbf{x}\right)\,.$
That is, $\mathcal{E}\subseteq\left\{h_{\boldsymbol{\theta}}\middle|h_{\boldsymbol{\theta}}\equiv h_{\boldsymbol{\theta}^{\star}}\right\}$ and therefore
$\mathbb{P}\left(\mathcal{E}\right)\leq\mathbb{P}\left(h_{\boldsymbol{\theta}}\equiv h_{\boldsymbol{\theta}^{\star}}\right)\,.$
Finally, to calculate the probability to sample $h_{\boldsymbol{\theta}}$ in $\mathcal{E}$ we count the number of constrained parameters - parameters which are either determined by $h_{\boldsymbol{\theta}^{\star}}$ or are 00 in $\mathcal{E}$. Looking at the dimensions of $\mathbf{W}_{11}^{\left(l\right)}$, $\mathbf{b}_{1}^{\left(l\right)}$, $\mathbf{b}_{2}^{\left(l\right)}$ and $\boldsymbol{\gamma}^{\left(l\right)}$, we deduce that there are exactly
@@
As in the proof of Theorem C.3, under the uniform prior over parameters $\mathcal{P}$,
@@
C.3 Convolutional Neural Network
@@
Definition C.7 (CNN restated).
For multi-channel inputs $\mathbf{x}^{j},\,\forall j\in\left[c_{0}\right]$, depth $L$, activation function $\sigma\left(\cdot\right)$, and multi-index channels number $D=\left(c_{1},\dots,c_{L},d_{s}\right)$, $\mathbf{h}^{\mathrm{CNN}}_{\boldsymbol{\theta}}$ is a convolutional network (CNN) defined recursively, starting with $\forall j\in\left[c_{0}\right]$, $\mathbf{f}_{\mathrm{CNN}}^{\left(0\right),j}\left(\mathbf{x}\right)=\mathbf{x}^{j}$, and then, for all $l\in\left[L\right]$ and $i\in\left[c_{l}\right]$, as
@@
$\mathbf{h}^{\mathrm{CNN}}_{\boldsymbol{\theta}}\left(\mathbf{x}\right)=$ $\operatorname{sign}\left(\mathrm{Vec}\left(\mathbf{f}_{\mathrm{CNN}}^{\left(L-1\right)}\left(\mathbf{x}\right)\right)^{\top}\mathbf{w}^{\left(L+1\right)}+b^{\left(L+1\right)}\right)\,,$
@@
$\mathbf{\boldsymbol{\theta}}=\left(\left\{\mathbf{K}^{\left(l\right)}\right\}_{l=1}^{L},\left\{\mathbf{b}^{\left(l\right)}\right\}_{l=1}^{L},\mathbf{w}^{\left(L+1\right)},b^{\left(L+1\right)}\right)\,,$
@@
Motivation. In order to tackle the analysis of a convolutional neural network at ease, we will define some new operations. We observe that our previous arguments from Section C.1 is agnostic to the convolution itself and relies only on parameter counting. In general, CNNs are analogous to FCNs in the sense that with spatial dimension of $1$, a convolutional layer is equivalent to a layer in a FCN, with each channel analogous to a neuron. Therefore, we define:
Definition C.9.
@@
and Def. C.9 we can write the multi-channel convolution in block form as
@@
Specifically, we can represent a convolutional layer, as defined in Def. C.7 by
@@
Theorem C.10 (Teacher equivalence probability for CNN restated).
For any activation function and any depth $L$, let $D^{\star}=\left(c_{1}^{\star},\dots,c_{L}^{\star},d_{s}^{\star}\right),\;D=\left(c_{1},\dots,c_{L},d_{s}\right)\in\mathbb{N}^{L+1}$ such that $D^{\star}\leq D$. If there exists some teacher $h_{\boldsymbol{\theta}^{\star}}\in\mathcal{H}^{\mathrm{CNN}}_{D^{\star}}$ and the prior is $\mathcal{P}=\mathcal{P}\left(\mathcal{H}^{\mathrm{CNN}}_{D}\right)$ then
@@
And by Lemma 3.2, $N=(\tilde{C}+3\log 2/\delta)/\varepsilon$ samples are enough to ensure that for posterior sampling (i.e. G&C), $\mathcal{L}(\mathcal{A}_{\mathcal{P}}(\mathcal{S}))\leq\varepsilon$ with probability $1-\delta$ over $\mathcal{S}\sim\mathcal{D}^{N}$ and the sampling.
@@
We follow the same strategy as in the proof of Theorem C.3. We define a sufficient condition on the parameters of the student network to ensure teacher equivalence, and then find the probability that this condition holds. Recall that for all $l\in\left[L\right]$, $\mathbf{K}^{\star\left(l\right)}$ and $\mathbf{b}^{\star\left(l\right)}$, where $\mathbf{K}_{i,j}^{\star\left(l\right)}\in\mathbb{R}^{k_{l}}$ is a convolution kernel with $k_{l}$ parameters, and $\mathbf{b}^{\star\left(l\right)}\in\mathbb{R}^{c_{l}}$, are the teacher’s $l^{\text{th}}$ layer’s convolution kernels and bias terms, respectively. Using the notation introduced in Def. C.9, the teacher’s convolution kernels can be arranged as
@@
=$\mathcal{E}=\left\{h_{\boldsymbol{\theta}}\in\mathcal{H}_{D}^{\text{CNN}}\middle|\begin{array}[]{l}\forall l=1,\dots,L\;\mathbf{K}^{\left(l\right),\left(11\right)}=\mathbf{K}^{\star\left(l\right)},\;\mathbf{K}^{\left(l\right),\left(12\right)}=\mathbf{0},\;\mathbf{b}_{1}^{\left(l\right)}=\mathbf{b}^{\star\left(l\right)},\\ \mathbf{w}^{\left(L+1\right),1}=\mathbf{w}^{\star\left(L+1\right)},\;\mathbf{w}^{\left(L+1\right),2}=\mathbf{0}_{d_{s}-d_{s}^{\star}},\;b^{\left(L+1\right)}=b^{\star\left(L+1\right)}\end{array}\right\}\,.$
We claim that any $h_{\boldsymbol{\theta}}\in\mathcal{E}$ is TE and can show this by induction over the layer $l$. With block notation, it is easy to see that this is identical to the proof of Theorem C.3 when substituting $\mathbf{W}^{\left(l\right)}$ with $\mathbf{K}^{\left(l\right)}$, $f_{\mathrm{FC}}^{\left(l\right)}$ with $\mathbf{f}_{\mathrm{CNN}}^{\left(l\right)}$, and standard matrix multiplication with convolution. Finally, we turn to finding $\mathbb{P}\left(\mathcal{E}\right)$. For a CNN $\mathbf{h}^{\mathrm{CNN}}_{\boldsymbol{\theta}}\in\mathcal{E}$, there are $c_{l}^{\star}\cdot c_{l-1}^{\star}$ convolution kernels set to equal the teacher’s, $c_{l}^{\star}\cdot\left(c_{l-1}-c_{l-1}^{\star}\right)$ kernels set to 00, and $c_{l}^{\star}$ bias terms. Each convolution kernel $\mathbf{K}_{i,j}^{\left(l\right)}$ is defined by $k_{l}$ parameters. From the linear layer, we need to account for $d_{s}$ weight parameters and one bias term. In total,
@@
C.4 Channel-Scaled Convolutional Neural Networks
Analogous to the FCN case, we present an additional CNN architecture. For simplicity, we state the definition of the architecture using the previously defined notation from C.9 and (40).
Definition C.11 (Channel Scaled CNN).
For multi-channel inputs $\mathbf{x}^{j},\,\forall j\in\left[c_{0}\right]$, depth $L$, activation function $\sigma\left(\cdot\right)$, and multi-index channels number $D=\left(c_{1},\dots,c_{L},d_{s}\right)$, $\mathbf{h}^{\mathrm{SCN}}_{\boldsymbol{\theta}}$ is a channel scaled convolutional neural network, or scaled convolutional network (SCN) for short, defined recursively, starting with $\mathbf{f}_{\mathrm{SCN}}^{\left(0\right)}\left(\mathbf{x}\right)=\mathbf{x}$, and then, for all $l\in\left[L\right]$, as
$\mathbf{f}_{\mathrm{SCN}}^{\left(l\right)}\left(\mathbf{x}\right)=$ $\left(\boldsymbol{\gamma}^{\left(l\right)}\odot\sigma\mathbf{K}^{\left(l\right)}*\mathbf{f}_{\mathrm{SCN}}^{\left(l-1\right)}+\mathbf{b}^{\left(l\right)}\right)$ (41)
$\mathbf{h}^{\mathrm{SCN}}_{\boldsymbol{\theta}}\left(\mathbf{x}\right)=$ $\operatorname{sign}\left(\mathrm{Vec}\left(\mathbf{f}_{\mathrm{SCN}}^{\left(L\right)}\left(\mathbf{x}\right)\right)^{\top}\mathbf{w}^{\left(L+1\right)}+b^{\left(L+1\right)}\right)\,,$